M11-4 Calculating Accounting Rate of Return, Payback Perlod [LO 11-1, 11-2 Blue Marlin Company is considering the purchase of new equipment for its factory. It will cost S250,000 and have a $50,000 salvage value in five years. The annual net income from the equipment is expected to be $30,000, and depreciation is $40,000 per year Calculate Blue Marlins annual rate of return and payback period for the equipment. (Do not round intermediate calculations. Round your Payback Period to 2 decimal places.) Annusl Rate ol Relurn Payback Period Years
Answer #1

Cost = $250,000

Annual Net Income = $30,000

Annual rate of Return = Annual Net Income/Investment

= 30,000/250,000 = 12%

Payback period is the time period in which initial investment is recovered.

Payback Period = Initial Investment/Annual Cash Flow

= 250,000/70,000

= 3.57 years

Working note: Annual Cash Flow = Annual Income + Depreciation Charge (since non-cash expenditure)

= $30,000+$40,000
= $70,000